The area of a triangle is 75 square meters. The altitude of the triangle is 5 meters longer than the base of the triangle. Find the base and the altitude of the triangle.
Answer by Alan3354(69443)

The area of a triangle is 75 square meters. The altitude of the triangle is 5 meters longer than the base of the triangle. Find the base and the altitude of the triangle.
---------
Area = b*h/2 = 75
b*h = 150
Find a pair of factors of 150 that differ by 5.